A body of mass m moves with velocity v and on the way meets (joins) another body of mass 2m, which was at rest.


 After merging the two bodies move rapidly:


 A) v2 = v / 3


 B) v2 = v / 2


 C) v2 = v / 4


 D) v2 = v


1
Expert's answer
2021-03-10T17:20:03-0500

According to the momentum conservation law, the momentum before the collision:


"p_1 = mv + 2m\cdot 0"

(since the second body did not move before the collsion) is equal to the momentum after the collision:


"p_2 =(m + 2m)v_2"

Thus, obtain:


"p_1 = p_2\\
mv = (m+2m)v_2\\
v_2 = \dfrac{v}{3}"

Answer. A.